OpenStudy (anonymous):

write an equation of a line that passes through (-4,3) and is perpendicular to the line y=7/5x-3/5

OpenStudy (anonymous):

these were the options A.5x+7y=1 B. 7x+5y=-13 C.7x+5y=-1 D.9x+5y=7

OpenStudy (anonymous):

Perpendicular to y=7/5x -3/5 is also perpendicular to y=7/5x. y=-5/7x + b is perpendicular to your line. That's because the a (7/5) of the first equation was opposed (times minus one) and inversed (denominators flipped) in the second equation. Now you can choose a parameter b so that y=-5/7x+b passes through (-4,3). Can you do that?

OpenStudy (anonymous):

denominator and numerator flipped, I meant

OpenStudy (anonymous):

so it wouldn't be A?